The equation e^(-Lx)=Lx has a root/zero between x=0 and x=1. This root is approximately at the midpoint of these two x-values. Find this root using a suitable algebra method.
L is 1.55
1
Expert's answer
2019-08-01T13:26:03-0400

Let Lx = y, equation e-y = y or yey = 1 have exact solution y = W(1), where W is Lambert W function, W(1) = Ω = Omega constant. Ω is transcendental number.

Recurrent relation for Ω is


"\varOmega_{n+1}=\frac{1+\varOmega_n}{1+e^{\varOmega_n}},\ \varOmega_1=1"

Using this recurrent sequence first 5 values for Ωi are: Ω1 = 1, Ω2 = 0.537883, Ω3 = 0.566987,

Ω4 = 0.567143, Ω5 = 0.567143.

Root of the equation f(y) = yey - 1 = 0 can be found using Newton's method.

"f'(y) = (ye^y-1)'=e^y+ye^y"

recurrent relation for Newton's method is


"y_{n+1}=y_n-\frac{f(y_n)}{f'(y_n)}=y_n-\frac{y_ne^{y_n}-1}{e^{y_n}+y_ne^{y_n}}, y_1=1"

Using this recurrent sequence first 5 values for y are y1 = 1, y2 = 0.68394, y3 = 0.577454,

y4 = 0.56723, y5 = 0.567143.

If L = 1.55, x = Ω/L = 0.567143/1.55 = 0.36589

Answer: If L = 1.55, x = 0.36589.